5. Lover Hardware Store had net credit sales of P 6,500,000 and cost of goods sold of P 5,500,000

for the year. The Accounts Receivable balances at the beginning and end of the year were P

600,000 and P 700,000, respectively. The Receivable turnover was:

a. 7.7 times

b. 10.8 times

c. 9.3 times

d. 10.0 times


1
Expert's answer
2022-03-08T15:46:01-0500

d. 10.0 times


Receivable turnover "=\frac{Net Credit Sales}{Average Accounts Receivable}"


"=\frac{P 6,500,000}{(P 600,000+P 700,000)÷2}=10"

Hence the correct answer is d.